# Measuring and Controlling Solution Degeneracy across Task-Trained Recurrent Neural Networks

### v2

@@ 3.1.1 Additional axes of task complexity
Changing memory demand. Of the four tasks, only Delayed Discrimination requires extended memory, as its performance depends on maintaining the first stimulus across a variable delay. See Appendix C for a quantification of each task’s memory demand. We increased the memory load in Delayed Discrimination by lengthening the delay period. This manipulation reduced degeneracy at the dynamical and behavioral levels but increased it at the weight level, mirroring the effect of increasing task dimensionality (Figure 4A).
Figure 4: Memory demand and auxiliary loss modulate degeneracy in distinct ways. In the Delayed Discrimination task, both manipulations reduce dynamical and behavioral degeneracy while increasing weight degeneracy. The auxiliary loss also induces additional line attractors in the network’s dynamics, as shown in (C).
Adding auxiliary loss. We next examined how adding an auxiliary loss affects solution degeneracy in the Delayed Discrimination task. Specifically, the network outputs both the sign and the magnitude of the difference between two stimulus values ($f_{2}-f_{1}$), using separate output channels for each. This manipulation added a second output channel and increased memory demand by requiring the network to track the magnitude of the difference between incoming stimuli. Consistent with our hypothesis, this manipulation reduced dynamical and behavioral degeneracy while increasing weight degeneracy (Figure 4B). Crucially, the auxiliary loss induced additional line attractors in the network dynamics, further structuring internal trajectories and aligning neural responses across networks (Figure 4C). While the auxiliary loss increases both output dimensionality and temporal memory demand, we interpret its effect holistically as a structured increase in task complexity.
@@ 3.2.1 Task complexity scales feature learning
Figure 5: More complex tasks drive stronger feature learning in RNNs. Increased input–output dimensionality leads to higher weight-change norms and lower kernel alignment. Error bars indicate ±$\pm 1$ standard error.
Neural networks can operate in either a lazy or rich learning regime [13, 82, 28, 7]. In the lazy regime, weights and internal features remain largely unchanged during training. In the rich (feature learning) regime, networks reshape their hidden representations and weights to capture task-specific structure [13, 29, 47, 82]. Harder tasks tend to push networks out of the lazy regime and into feature learning, leading to stronger adaptation of weights and internal representations [8, 46].
@@
We hypothesize that the increased weight degeneracy observed in harder tasks reflects stronger feature learning. Specifically, harder tasks may force network weights to travel farther from their initialization. If more complex task variants, like those in Section 3.1, truly induce greater feature learning, then networks should traverse a greater distance in weight space, resulting in more dispersed final weights. To test this idea, we measured feature learning strength in networks trained on different task variants using two complementary metrics  [53, 29]: Weight-change norm: $\left\|\mathbf{W}_{T}-\mathbf{W}_{0}\right\|_{F}$, where larger values indicate stronger feature learning. Kernel alignment (KA): measures the directional change of the neural tangent kernel (NTK) before and after training: $\operatorname{KA}\bigl{(}K^{(f)},K^{(0)}\bigr{)}=\frac{\operatorname{Tr}\left(K^{(f)}K^{(0)}\right)}{\left\|K^{(f)}\right\|_{F}\left\|K^{(0)}\right\|_{F}}$, where $K=\nabla_{W}\hat{y}^{\top}\nabla_{W}\hat{y}$. Lower KA indicates greater NTK rotation and thus stronger feature learning.
@@
More complex tasks consistently drive stronger feature learning and greater dispersion in weight space, as reflected by increasing weight-change norm and decreasing kernel alignment across all tasks (Figure 5).
3.3 Controlling feature learning reshapes degeneracy across levels
@@
Our earlier results show that harder tasks induce stronger feature learning, which in turn shapes the dispersion of solutions in the weight space. To test whether feature learning causally affects degeneracy across weight, dynamics, and behavior, we used a principled parameterization known as maximum update parameterization ($\mu P$), which enables stable feature learning across network widths—even in the infinite-width limit [7, 13, 28, 82]. In this setup, a single hyperparameter—$\gamma$—controls the strength of feature learning: higher $\gamma$ values induce a richer feature-learning regime. Under this parameterization, the network update rule, initialization, and learning rate are scaled with respect to network width $N$. For the Adam optimizer, the output is scaled as $f(t)=\frac{1}{\gamma N}W_{\text{readout}}\phi(h(t))$. The hidden state update is scaled as $h(t+1)-h(t)=\tau\left(-h(t)+\frac{1}{N}J\phi(h(t))+Ux(t)\right)$, where $J_{ij}\sim\mathcal{N}(0,N)$. In this regime, the learning rate scales as $\eta=\gamma\eta_{0}$. A detailed explanation of $\mu P$ and its relationship to the standard parameterization is in Appendix  G and H. For each task, we trained networks with multiple $\gamma$ values and confirmed that larger $\gamma$ consistently induces stronger feature learning, as evidenced by increased weight-change norm and decreased kernel alignment (Appendix I).
@@
Figure 6: Stronger feature learning reduces dynamical degeneracy but increases weight and behavioral degeneracy. Panels show degeneracy at the dynamical, weight, and behavioral levels (top to bottom). Shaded area indicates ±$\pm 1$ standard error.
@@
We observed that stronger feature learning reduced degeneracy at the dynamical level but increased it at the weight level. Intuitively, when $\gamma$ is high, networks tend to learn similar task-specific features and converge to consistent dynamics and behavior. In contrast, lazy networks (with small $\gamma$) rely on their initial random features, leading to more divergent solutions across seeds—even though their weights move less overall (Figure 6). This finding aligns with prior work in feedforward networks, where feature learning was shown to reduce the variance of the neural tangent kernel across converged models [8]. At the behavioral level, however, increasing feature-learning strength leads networks to overfit the training distribution (Appendix F.2). We hypothesize that stronger feature learning exacerbates overfitting, increasing both average OOD loss and the variability of OOD behavior across models (Figure 6[4, 20, 38, 51]. Although stronger feature learning increases behavioral degeneracy, this may partially reflect overfitting to the training distribution rather than inherent task difficulty—an effect we highlight in Appendix F.2. Clarifying the mechanistic link between dynamical and behavioral degeneracy remains an important direction for future work.
@@ 3.4 Larger networks yield more consistent solutions across levels
Figure 7: Larger networks reduce degeneracy across weight, dynamics, and behavior. Controlling for feature learning strength, wider RNNs yield more consistent solutions across all three levels of analysis. Panels show degeneracy at the dynamical, weight, and behavioral levels (top to bottom). Shaded area indicates ±$\pm 1$ standard error.
Prior work in machine learning and optimization shows that over-parameterization improves convergence by helping gradient methods escape saddle points [41, 64, 19, 2, 86]. We therefore hypothesized that larger RNNs would converge to more consistent solutions across seeds. However, increasing width also pushes models toward the lazy neural tangent kernel (NTK) regime, where feature learning is suppressed [39, 47, 13, 82, 28]. To disentangle these competing effects, we again use the $\mu P$ parameterization, which holds feature learning strength constant (via fixed $\gamma$) while scaling width. Although larger networks may yield more consistent solutions via self-averaging, this outcome is not guaranteed without controlling for feature learning. In standard RNNs, increasing width often induces lazier dynamics, which can paradoxically increase dynamical degeneracy rather than reduce it. The $\mu P$ setup enables us to isolate the size effect cleanly.
Across all tasks, larger networks consistently exhibit lower degeneracy at the weight, dynamical, and behavioral levels, producing more consistent solutions across random seeds (Figure 7). This pattern aligns with findings in vision and language models, where wider networks converge to more similar internal representations [61, 44, 71, 81, 43]. In recurrent networks, only a few studies have investigated this “convergence-with-scale” effect using representation-based metrics [61, 65]. Our results extend these findings by (1) focusing on neural computations across time (i.e., neural dynamics) rather than static representations, and (2) demonstrating convergence-with-scale across weight, dynamical, and behavioral levels in RNNs.
@@ 3.5 Structural regularization reduces solution degeneracy
Figure 8: Low-rank and sparsity regularization reduce solution degeneracy across all levels. On the Delayed Discrimination task, both regularizers lower degeneracy in dynamics, weights, and behavior. Shaded area indicates ±$\pm 1$ standard error.
Low-rank and sparsity constraints are widely used structural regularizers in neuroscience-inspired modeling and efficient machine learning [57, 6, 66, 36, 32]. A low-rank penalty compresses the weight matrices into a few dominant modes, while an $\ell_{1}$ penalty drives many parameters to zero and induces sparsity. In both cases, task-irrelevant features are pruned, nudging independently initialized networks toward more consistent solutions on the same task. To test this idea, we augmented the task loss with either a nuclear-norm penalty on the recurrent weights $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{\text{rank}}\sum_{i=1}^{r}\sigma_{i}$, where $\sigma_{i}$ are the singular values of the recurrent matrix, or an $\ell_{1}$ sparsity penalty: $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{\ell_{1}}\sum_{i}\left|w_{i}\right|$. We focused on the Delayed Discrimination task to control for baseline difficulty, but both regularizers consistently reduced degeneracy across all levels—and similar effects hold in other tasks (Appendix K, Figure 8).
@@ 4 Discussion
Higher Task complexity (contravariant) ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{1,0.23,0.13}\boldsymbol{\uparrow}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$
More Feature learning (contravariant) ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{1,0.23,0.13}\boldsymbol{\uparrow}}$ ${\color[rgb]{1,0.23,0.13}\boldsymbol{\uparrow}}$
Larger Network size (covariant) ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$
Regularization (covariant) ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$ ${\color[rgb]{0,0.5,1}\boldsymbol{\downarrow}}$
@@
Two consistent patterns emerged from this analysis. First, increasing task complexity or boosting feature learning produced a contravariant effect: dynamical degeneracy decreased while weight degeneracy increased. Second, increasing network size or applying structural regularization reduced degeneracy at both the weight and dynamical levels—that is, a covariant effect. Here, covariant and contravariant refer to the relationship between weight and dynamic degeneracy—not whether degeneracy increases or decreases overall. For example, task complexity and feature learning reduce dynamical degeneracy but increase weight degeneracy, whereas network size and regularization reduce both.
@@
We also observed that the relationship between dynamical and behavioral degeneracy depends on the varying factor. For instance, stronger feature learning leads to more consistent neural dynamics on the training task but greater variability in OOD generalization This suggests that tightly constrained dynamics on the training set do not guarantee more consistent behavior on OOD inputs. This highlights the need for further empirical and theoretical work on how generalization depends on the internal structure of task-trained networks [50, 14, 75]. This divergence highlights a key open question: how much of behavioral consistency generalizes beyond training-aligned dynamics, and what task or network factors drive this decoupling?
@@
These knobs allow researchers to tune the level of degeneracy in task-trained RNNs to suit specific research questions or application needs. For example, researchers may want to suppress degeneracy to study common mechanisms underlying a neural computation. Conversely, to probe individual differences, they can increase degeneracy to expose solution diversity across independently trained networks [9, 62, 85]. Our framework also supports ensemble-based modeling of brain data. By comparing dynamical and behavioral degeneracy across trained networks, it may be possible to match inter-individual variability in models to that observed in animals—helping capture the full distribution of task-solving strategies [63, 3, 11, 68].
@@
Although our analyses use artificial networks, several of the mechanisms we uncover may translate directly to experimental neuroscience. For example, introducing an auxiliary sub-task during behavioral shaping—mirroring our auxiliary-loss manipulation—could constrain the solution space animals explore, thereby reducing behavioral degeneracy [37]. Finally, our contrasting findings motivate theoretical analysis—e.g., using linear RNNs—to understand why some factors induce contravariant versus covariant relationships across behavioral, dynamical, and weight-level degeneracy.
In summary, our work takes a first step toward addressing two classic puzzles in task-driven modeling: (1) What factors shape the variability across independently trained networks? (2) To what extent can we trust interpretations drawn from a single model?
@@
We present a unified framework for quantifying solution degeneracy in task-trained RNNs, identify the key factors that shape the solution landscape, and provide practical guidance for controlling degeneracy to match specific research goals in neuroscience and machine learning.
@@
Limitations and future directions. This work considers networks equivalent if they achieve similar training loss. Future work could extend the framework to tasks with multiple qualitatively distinct solutions, to examine whether specific factors bias the distribution of networks across those solutions. Another open question is the observed decoupling between dynamical and behavioral degeneracy: how much of behavioral consistency generalizes beyond training-aligned dynamics, and what task or network factors drive this divergence?

### v3

@@ 3.1.1 Additional axes of task complexity
Figure 4: Increasing memory demand or adding auxiliary loss changes task complexity, which in turn modulates degeneracy. In the Delayed Discrimination task, both manipulations reduce dynamical and behavioral degeneracy [temporal generalization] while increasing weight degeneracy. The auxiliary loss also induces additional line attractors in the network’s dynamics, as shown in (C).
@@
Changing memory demand. Of the four tasks, only Delayed Discrimination requires extended memory, as its performance depends on maintaining the first stimulus across a variable delay. See Appendix D for a quantification of each task’s memory demand. We increased the memory load in Delayed Discrimination by lengthening the delay period. This manipulation reduced degeneracy at the dynamical and behavioral levels but increased it at the weight level, mirroring the effect of increasing task dimensionality (Figure 4A).
Adding auxiliary loss. We next examined how adding an auxiliary loss affects solution degeneracy in the Delayed Discrimination task. Specifically, the network outputs both the sign and the magnitude of the difference between two stimulus values ($f_{2}-f_{1}$), using separate output channels for each. This manipulation added a second output channel and increased memory demand by requiring the network to track the magnitude of the difference between incoming stimuli. Consistent with our hypothesis, this manipulation reduced dynamical and behavioral degeneracy [temporal generalization] while increasing weight degeneracy (Figure 4B). Crucially, the auxiliary loss induced additional line attractors in the network dynamics, further structuring internal trajectories and aligning neural responses across networks (Figure 4C). While the auxiliary loss increases both output dimensionality and temporal memory demand, we interpret its effect holistically as a structured increase in task complexity.
@@ 3.2.1 Task complexity scales feature learning
In deep learning theory, neural networks can either solve tasks using their random features at initialization, or adapt their weights and internal features to capture task specific structure [chizat2019lazy, woodworth2020kernel, geiger2020disentangling, bordelon_self-consistent_2022]. These are referred to as the lazy learning regime, where weights and internal features remain largely unchanged during training, and the rich learning, or feature learning regime, where networks reshape their hidden representations and weights to capture task-specific structure [chizat2019lazy, george2022lazy, lee2019wide, woodworth2020kernel]. As the complexity of a task grows, the initial random features no longer suffice to solve it, pushing the network beyond the lazy regime and into feature learning, where weights and internal representations adapt more substantially. [bordelon_dynamics_2024, kumar2023grokking]. If more complex task variants, like those in Section 3.1, truly induce greater feature learning, then networks should adapt more from their initializations and traverse a greater distance in the weight space, resulting in more dispersed final weights.
@@
Figure 5: More complex tasks drive stronger feature learning in RNNs. Increased input–output dimensionality leads to higher weight-change norms ($\Delta W||_{F}$) and lower kernel alignment (KA). Error bars indicate $\pm 1$ standard error.
@@
We therefore hypothesize that the increased weight degeneracy observed in harder tasks reflects stronger feature learning within the network. To test this idea, we measured feature learning strength in networks trained on different task variants using two complementary metrics  [liu2023connectivity, george2022lazy]: Weight-change norm: $\left\|\mathbf{W}_{T}-\mathbf{W}_{0}\right\|_{F}$, where larger values indicate stronger feature learning. Kernel alignment (KA): The geometry of learning under gradient descent can be described by the neural tangent kernel (NTK), which captures how weight updates affect the network outputs. The NTK is defined by $K=\nabla_{W}\hat{y}^{\top}\nabla_{W}\hat{y}$ where $\hat{y}$ denotes the network output. KA measures the directional change of the NTK before and after training: $\operatorname{KA}\bigl(K^{(T)},K^{(0)}\bigr)=\tfrac{\operatorname{Tr}\left(K^{(T)}K^{(0)}\right)}{\left\|K^{(T)}\right\|_{F}\left\|K^{(0)}\right\|_{F}}$. Lower KA indicates greater NTK rotation and thus stronger feature learning.
@@
We find that more complex tasks consistently drive stronger feature learning and greater dispersion in weight space, as reflected by increasing weight-change norm and decreasing kernel alignment across all tasks (Figure 5).
3.2.2 Controlling feature learning reshapes degeneracy across levels
@@
Our earlier results show that harder tasks induce stronger feature learning, which in turn shapes the dispersion of solutions in the weight space. To test whether feature learning causally affects degeneracy, we used a principled network parameterization known as maximum update parameterization ($\mu P$), which allows stable feature learning across network widths, even in the infinite-width limit [bordelon_self-consistent_2022, chizat2019lazy, geiger2020disentangling, woodworth2020kernel]. In this setup, a single hyperparameter ($\gamma$) controls the strength of feature learning: higher $\gamma$ values induce a richer feature-learning regime. Under this parameterization, the network update rule, initialization, and learning rate are scaled with respect to network width $N$. For the Adam optimizer, the output is scaled as $f(t)=\frac{1}{\gamma N}W_{\text{readout}}\phi(h(t))$. The hidden state update is scaled as $h(t+1)-h(t)=\tau\left(-h(t)+\frac{1}{N}J\phi(h(t))+Ux(t)\right)$, where $J_{ij}\sim\mathcal{N}(0,N)$ are the recurrent weights and $\phi$ is the tanh nonlinearity. The learning rate scales as $\eta=\gamma\eta_{0}$. A detailed explanation of $\mu P$ and its relationship to the standard parameterization is in Appendix  K and L. For each task, we trained networks with multiple $\gamma$ values and confirmed that larger $\gamma$ consistently induces stronger feature learning, as evidenced by increased weight-change norm and decreased kernel alignment (Appendix M).
@@
Figure 6: Stronger feature learning reduces dynamical degeneracy but increases weight and behavioral degeneracy. Panels show degeneracy at the dynamical, weight, and behavioral levels (top to bottom). Shaded area indicates $\pm 1$ standard error.
@@
We observed that stronger feature learning reduced degeneracy at the dynamical level but increased it at the weight level. We see that when $\gamma$ is high, networks tend to learn similar task‑specific features and converge to consistent dynamics and behavior. In contrast, lazy networks (with small $\gamma$) rely on their initial random features, leading to more divergent solutions across seeds—even though their weights move less overall (Figure 6). This finding aligns with prior work in feedforward networks, where feature learning was shown to reduce the variance of the neural tangent kernel across converged models [bordelon_dynamics_2024]. At the behavioral level, however, increasing feature‑learning strength leads networks to overfit the training distribution (Appendix J.2). We hypothesize that stronger feature learning exacerbates overfitting, increasing both average OOD loss and the variability of OOD behavior across models (Figure 6[bansal2021revisiting, duan2020unsupervised, huh2024platonic, li2016convergent]. Although stronger feature learning increases behavioral degeneracy [temporal generalization], this may partially reflect overfitting to the training distribution, an effect we highlight in Appendix J.2. Clarifying the mechanistic link between dynamical and behavioral degeneracy [temporal generalization] remains an important direction for future work. In Appendix I, we demonstrate that the observed effects of feature learning on degeneracy both interpolates smoothly within the range of $\gamma$ values and extrapolates beyond the range reported in Figure 6.
@@ 3.3 Larger networks yield more consistent solutions across levels
Figure 7: Larger networks reduce degeneracy across weight, dynamics, and behavior. After controlling for feature learning strength ($\gamma=1$ held constant across network widths), wider RNNs yield more consistent solutions across all three levels of analysis. Panels show degeneracy at the dynamical, weight, and behavioral levels (top to bottom). Shaded area indicates $\pm 1$ standard error.
Prior work in machine learning and optimization shows that over-parameterization improves convergence by helping gradient methods escape saddle points [kawaguchi2016deep, nguyen2017loss, du2018gradient, allen2019convergence, zou2018sgd, simsek2021geometrylosslandscapeoverparameterized, martinelli2024expandandclusterparameterrecoveryneural]. We therefore hypothesized that larger RNNs would converge to more consistent solutions across seeds. However, increasing width also tends to push models towards the lazy regime, where feature learning is suppressed [jacot2018neural, lee2019wide, chizat2019lazy, woodworth2020kernel, geiger2020disentangling]. To disentangle these competing effects, we again use the $\mu P$ parameterization, which holds feature learning strength constant (via fixed $\gamma$) while scaling width. Although larger networks may yield more consistent solutions via self-averaging, this outcome is not guaranteed without controlling for feature learning. In standard RNNs, increasing width often induces lazier dynamics, which can paradoxically increase dynamical degeneracy rather than reduce it. The $\mu P$ setup enables us to isolate the size effect cleanly.
Across all tasks, larger networks consistently exhibit lower degeneracy at the weight, dynamical, and behavioral levels, producing more consistent solutions across random seeds (Figure 7). Our dense sweep over 12 intermediate network sizes from 32 to 512 on the 3-Bits Flip Flop task in Appendix I further confirms the observed effect of network width on degeneracy. This pattern aligns with findings in vision and language models, where wider networks converge to more similar internal representations [morcos2018pwcca, kornblith2019cka, raghu2017svcca, wolfDynamicalModelsCortical2014, contrasim2024, huh2024platonic]. In recurrent networks, only a few studies have investigated this “convergence-with-scale” effect using representation-based metrics [morcos2018pwcca, nguyen2021wide]. Our results extend these findings by (1) focusing on neural computations across time (i.e., neural dynamics) rather than static representations, and (2) demonstrating convergence-with-scale across weight, dynamical, and behavioral levels in RNNs.
@@ 3.4 Structural regularization reduces solution degeneracy
Figure 8: Low-rank and sparsity regularization reduce solution degeneracy across all levels. On the Delayed Discrimination task, both regularizers lower degeneracy in dynamics, weights, and behavior. Shaded area indicates $\pm 1$ standard error.
Low-rank and sparsity constraints are widely used structural regularizers in neuroscience-inspired modeling and efficient machine learning [mastrogiuseppe2018linking, beiran2020shaping, olshausen1996emergence, han2015learning, glorot2011deep]. A low-rank penalty compresses the weight matrices into a few dominant modes, while an $\ell_{1}$ penalty drives many parameters to zero and induces sparsity. In both cases, task-irrelevant features are pruned, nudging independently initialized networks toward more consistent solutions on the same task. To test this idea, we augmented the task loss with either a nuclear-norm penalty on the recurrent weights $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{\text{rank}}\sum_{i=1}^{r}\sigma_{i}$, where $\sigma_{i}$ are the singular values of the recurrent matrix, or an $\ell_{1}$ sparsity penalty: $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{\ell_{1}}\sum_{i}\left|w_{i}\right|$. We focused on the Delayed Discrimination task to control for baseline difficulty, and observe that both regularizers consistently reduced degeneracy across all levels. Similar effects hold in other tasks (Appendix O, Figure 8) and intermediate regularization strengths (Appendix I ).
@@ 4 Discussion
[…]
@@
Two consistent patterns emerged from this analysis. First, increasing task complexity or boosting feature learning produced a contravariant effect: dynamical degeneracy decreased while weight degeneracy increased. Second, increasing network size or applying structural regularization reduced degeneracy at both the weight and dynamical levels—that is, a covariant effect. Here, covariant and contravariant refer to the relationship between weight and dynamic degeneracy, not whether degeneracy increases or decreases overall. For example, task complexity and feature learning reduce dynamical degeneracy but increase weight degeneracy, whereas network size and regularization reduce both.
@@
We also observed that the relationship between dynamical and behavioral degeneracy depends on the varying factor. For instance, stronger feature learning leads to more consistent neural dynamics on the training task but greater variability in OOD generalization This suggests that tightly constrained dynamics on the training set do not guarantee more consistent behavior on OOD inputs. This highlights the need for further empirical and theoretical work on how generalization depends on the internal structure of task-trained networks [li2024, Cohen2020, Sorscher2022]. This divergence highlights a key open question: how much of behavioral consistency generalizes beyond training-aligned dynamics, and what task or network factors drive this decoupling?
@@
These knobs allow researchers to tune the level of degeneracy in task-trained RNNs to suit specific research questions or application needs. For example, researchers may want to suppress degeneracy to study common mechanisms underlying a neural computation. Conversely, to probe individual differences, they can increase degeneracy to expose solution diversity across independently trained networks [bouthillier2021accounting, morik2005sloppy, yang2022does, singh2023emergent]. Our framework also supports ensemble-based modeling of brain data. By comparing dynamical and behavioral degeneracy across trained networks, it may be possible to match inter-individual variability in models to that observed in animals, helping capture the full distribution of task-solving strategies [NegronOyarzo2018, Ashwood2022, Cazettes2023, Pagan2025].
@@
Although our analyses use artificial networks, several of the mechanisms we uncover may translate directly to experimental neuroscience. For example, introducing an auxiliary sub-task during behavioral shaping, which mirrors our auxiliary-loss manipulation, could constrain the solution space animals explore, thereby reducing behavioral degeneracy [howard2002control]. Finally, our contrasting findings motivate theoretical analysis, e.g., using linear RNNs to understand why some factors induce contravariant versus covariant relationships across behavioral, dynamical, and weight-level degeneracy.
@@
In summary, our work takes a first step toward addressing this classic puzzle in task-driven modeling: What factors shape the variability across independently trained networks? We present a unified framework for quantifying solution degeneracy in task-trained RNNs, identify the key factors that shape the solution landscape, and provide practical guidance for controlling degeneracy to match specific research goals in neuroscience and machine learning.
@@
Limitations and future directions. This work considers networks equivalent if they achieve similar training loss. Future work could extend the framework to tasks with multiple qualitatively distinct solutions, to examine whether specific factors bias the distribution of networks across those solutions. Another open question is the observed decoupling between dynamical and behavioral degeneracy: how much of behavioral consistency generalizes beyond training-aligned dynamics, and what task or network factors drive this divergence.